OpenStudy (anonymous):

What would be the difference quotient ( ill insert the equation) for the given function? (f(x) = a, where x is the variable, h is the change in the variable, and a is a constant

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h} \]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

the derivative of a constant is 0. the difference quotient is 1. try evaluating it (x + h - x)/h = h/h = 1

OpenStudy (anonymous):

@bloopman thank you <3 xoxoxo you are a life saver hun

OpenStudy (skullpatrol):

did you learn anything?